OpenStudy (anonymous):

Prove that a line that divides two sides of a triangle proportionally is parallel to the third side. Be sure to create and name the appropriate geometric figures.

OpenStudy (anonymous):

@ranga Thanks again

OpenStudy (ranga):



OpenStudy (ranga):

Statement Reason DB/AD = EC/AE Given (Line divides 2 sides of a triangle proportionally) DB/AD+1 = EC/AE+1 Addition Property of Equality or (DB+AD)/AD = (EC+AE)/AE or AB/AD = AC/AE Included angle A = included Reflexive Property angle A Triangles ADE and ABC are SAS for Similarity Theorem. similar Angle ADE = Angle ABC Corresponding angles of similar triangles are congruent. DE is parallel BC Corresponding angles ADE and ABC are congruent.